OpenStudy (tiffanymak1996):

Let f(x,y) = (x^2 +y^2)^(2/3). Show that fx(x,y) = {4x/((3)(x^2+y^2)^1/3), (x,y) does not equal (0,0) and 0, (x,y)=(0,0)

OpenStudy (tiffanymak1996):

@zepdrix help?

OpenStudy (tiffanymak1996):

I can do the first part, but not the second part.

OpenStudy (tiffanymak1996):

this is partial differentiation again, thanks for helping.

zepdrix (zepdrix):

So you were able to find \(\large f_x(x,y)\) without any trouble? I don't understand what the next part is saying D: hmm

OpenStudy (tiffanymak1996):

yes, but for the next part, I need to show that fx is 0 when x=0 and y=0.

OpenStudy (tiffanymak1996):

If i differentiate it as usual, i'll end up with (x^2 +y^2)^1/3 at the bottom, and a denominator can't be zero, so...

zepdrix (zepdrix):

Oh hmmm D: Is there some type of limiting process we can do? Hmmmm

OpenStudy (tiffanymak1996):

can we have \[\lim_{(x,y) \rightarrow (0,0)} \frac{ 4x }{ 3(x^2+y^2)^\frac{ 1 }{ 3 } }\]

OpenStudy (tiffanymak1996):

since (x,y) can't be (0,0), instead of substituting it, we can limit it to (0,0)? I'm not sure.

zepdrix (zepdrix):

Hmm yah maybe that would work. I was thinking maybe we use the limit definition of the derivative. But I dunno :d \[\Large f_x(x,y)=\lim_{h\to0}\frac{\left[(x+h)^2+y^2\right]^{2/3}-\left[x^2+y^2\right]^{2/3}}{h}\] \[\Large f_x(0,0)=\lim_{h\to0}\frac{\left[(0+h)^2+0^2\right]^{2/3}-\left[0^2+0^2\right]^{2/3}}{h}\]

OpenStudy (tiffanymak1996):

yeah, that works, nice, thanks!

zepdrix (zepdrix):

Oh cool! ^^